# Fair Generation without Unfair Distortions: Debiasing Text-to-Image Generation with Entanglement-Free Attention

### v1

@@ Abstract
Recent advancements in diffusion-based text-to-image (T2I) models have enabled the generation of high-quality and photorealistic images from text descriptions. However, they often exhibit societal biases related to gender, race, and socioeconomic status, thereby reinforcing harmful stereotypes and shaping public perception in unintended ways. While existing bias mitigation methods demonstrate effectiveness, they often encounter attribute entanglement, where adjustments to attributes relevant to the bias (i.e., target attributes) unintentionally alter attributes unassociated with the bias (i.e., non-target attributes), causing undesirable distribution shifts. To address this challenge, we introduce Entanglement-Free Attention (EFA), a method that accurately incorporates target attributes (e.g., White, Black, Asian, and Indian) while preserving non-target attributes (e.g., background details) during bias mitigation. At inference time, EFA randomly samples a target attribute with equal probability and adjusts the cross-attention in selected layers to incorporate the sampled attribute, achieving a fair distribution of target attributes. Extensive experiments demonstrate that EFA outperforms existing methods in mitigating bias while preserving non-target attributes, thereby maintaining the output distribution and generation capability of the original model.
Figure 1: Generated images from the pretrained model (e.g., Stable Diffusion [1]), previous debiasing methods (e.g., UCE [10] and Interpret Diffusion [14]), and ours. Previous debiasing methods often alter the non-target attributes (e.g., house) while modifying bias-related attributes (e.g., racial attributes). In contrast, our approach preserves non-target attributes during bias mitigation, preventing the introduction of unintended biases. The corresponding aligned images were generated using the same random seed.
* indicates equal contribution.
@@ 1 Introduction
To address this issue, we propose Entanglement-Free Attention (EFA), a method that accurately incorporates target attributes while preserving non-target attributes. In this work, we focus on human-centric bias, which plays a crucial role in shaping societal perceptions. EFA promotes fair representation of predefined target attributes (e.g., White, Black, Asian, and Indian) by randomly sampling one at inference time to ensure balanced outputs. To guide the model to focus on semantically relevant regions (e.g., face and body), EFA is trained to enhance target attributes that exhibit high semantic relevance to the counterfactual prompts, while human segmentation masks further constrain its effect to human regions. By employing both target and counterfactual prompts in training, EFA learns to dynamically adjust the enhancement strength based on the presence of the target attribute, applying stronger modifications when it is underrepresented. At inference time, EFA operates without masks by internally localizing bias-relevant regions, thereby effectively disentangling target and non-target attributes while preserving the original model’s generative capacity and contextual consistency.
@@ 2 Related works
Figure 2: Overview of our method. (a): EFA is applied to selected layers and trained to reconstruct the image $\mathbf{x}_{0}$ (a-1) when given $p_{a}$, the prompt containing the target attribute $a$, or (a-2) its counterfactual prompt, $p_{a}^{\text{cf}}$. We leverage human segmentation mask $M$ to minimize the effect of EFA on the attention map $A$ of non-target attribute regions. (b): During the inference, our method generates outputs with a fair distribution of target attributes using the EFA of the randomly sampled target attribute.
@@
Existing studies on mitigating bias in T2I generative models have mainly followed two approaches. The first approach involves directly modifying the parameters of diffusion-based T2I models. For instance, Orgad et al. [18] and Gandikota et al. [10] proposed a method to adjust the weights of the cross-attention module to alleviate the bias. Similarly, Shen et al. [25] introduced a method that finetunes the text encoder to alleviate face-related biases such as gender and race. However, since finetuning is conducted on a significantly smaller dataset compared to the pretraining dataset, such approach may potentially degrade original output quality or diversity. Furthermore, the approach of Shen et al. [25] requires substantial computational costs for training the backbone model.
@@
Another line of research has explored inference-time steering approaches, where the pretrained model remains unchanged while adjustments are applied during inference. Friedrich et al. [9] proposed a bias mitigation framework that employs semantic guidance [4] to manipulate the target attributes by referring to registered information of a concept (e.g., nurse) and its bias (e.g., gender). However, this cannot mitigate bias for unregistered concepts, limiting its practicality. Chuang et al. [7] have developed a method to remove bias from text embeddings, and Kim et al. [12] trains the soft prompts for mitigating bias. Beyond employing the text embeddings, Shrestha et al. [26] introduced a technique that projects a reference image into text-compatible tokens, allowing the model to leverage the features of external images, such as gender or race for image generation. Additionally, Parihar et al. [19] and Li et al. [14] proposed methods to manipulate the h-space [13] of the model during inference. While these methods aim to mitigate bias without directly altering the backbone parameters, they struggle to preserve the original context of the generated images.
@@
Recently, Zhou et al. [30] addressed association-engendered stereotypes, which emerge from co-occurrence of multiple objects in an image. They rely on a CLIP model trained to predict biases from prompts, requiring prior knowledge of the biases to be mitigated. In contrast, our approach aims to address subtle biases induced by contextual variations beyond the predefined bias definitions. For instance, if a particular social group is predominantly depicted in low-resource work environments, while another group appears in corporate office settings. These context variations can reinforce cognitive biases, subtly implying disparities in professional status. Furthermore, it is impractical to anticipate all possible forms of such bias in advance.
@@ 3.3 Training
Dataset construction. To train $\text{EFA}_{C}^{a_{i}}$ to accurately represent a target attribute $a_{i}$ (e.g., female) while preserving non-target attributes, we create a dataset $\mathcal{D}_{a_{i}}$ that captures $a_{i}$ and provides explicit supervision on the human regions that contain $a_{i}$. The dataset is constructed as follows. First, we generate images using the original model with prompts explicitly describing $a_{i}$ (e.g., “a female person”). For each image, we extract a human segmentation mask $M$ using Grounded SAM2 [21, 16], assigning 1 to human regions and 0 elsewhere. Since this study focuses on human-centric bias, we utilize a human segmentation mask to serve as the basis for regulating the influence of $a_{i}$ during training. The resulting $\mathcal{D}_{a_{i}}$ consists of pairs of images and their segmentation masks.
Training EFA. At inference time, the extent of the features that need to be adjusted by EFA to faithfully reflect $a_{i}$ may vary across inputs. For instance, with the prompt “nurse”, intermediate features may exhibit a high level of femininity, requiring minimal adjustment by $\text{EFA}_{\text{gender}}^{\text{female}}$. In contrast, intermediate features of the prompt “doctor” may have stronger masculinity, requiring enhanced representation of female by $\text{EFA}_{\text{gender}}^{\text{female}}$ to incorporate ‘female’ into the output. To address such variability, $\text{EFA}_{C}^{a_{i}}$ is trained in two different scenarios : (1) when intermediate features contain the attribute $a_{i}$ (e.g., female), and (2) when intermediate features contain its counterfactual attribute $a_{i}^{\text{cf}}\in\mathcal{A}_{C}\setminus\{a_{i}\}$ (e.g., male). We denote $a_{i}$ as $a$ throughout this section for notational simplicity.
@@
$\mathcal{L}_{\text{reg}}={\lVert E(\text{softmax}([\frac{QK^{\top}}{\sqrt{d}},\text{AP}_{a}(\mathbf{z}_{t})]),V_{a})\rVert_{1}},$
where $\mathbf{z}_{t}=\phi(\textbf{x}_{t},p_{a})$, and the function $E(A,V)$ extracts the subset of attention values in the attention map $A$ corresponding to $V$. For simplicity, we denote $E(\text{softmax}([\frac{QK^{\top}}{\sqrt{d}},\text{AP}_{a}(\mathbf{z}_{t})]),V_{a})$ as $A_{a}$ in Fig. 2. This regularization loss minimizes unnecessary modifications introduced by $\text{EFA}_{C}^{a}$, ensuring that the model preserves the original cross-attention output.
@@
where $\epsilon_{\theta}$ denotes the U-Net with $\text{EFA}_{C}^{a}$, and $\odot$ represents the element-wise multiplication. Through the reconstruction process, $\text{EFA}_{C}^{a}$ is encouraged to enhance $a$ in regions where cross-attention assigns high similarity to $a^{\text{cf}}$, reflecting their semantic relevance. Using human masks ensures that $a$ is accurately reconstructed in human regions while mitigating the influence of $a^{\text{cf}}$.
Although $\mathcal{L}_{\text{reg}}$ reduces unintended modifications, it does not explicitly confine its impact on non-human regions, potentially affecting non-target attributes. To address this, we impose an additional L1 regularization on the attention values corresponding to $V_{a}$ as follows:
$\mathcal{L}_{\text{reg}}^{\text{cf}}={\lVert(1-\bar{M})\odot E(\text{softmax}([\frac{QK^{\top}}{\sqrt{d}},\text{AP}_{a}(\mathbf{z}_{t}^{\text{cf}})]),V_{a})\rVert_{1}},$
where $\mathbf{z}_{t}^{\text{cf}}=\phi(\textbf{x}_{t},p_{a}^{\text{cf}})$, and $\bar{M}$ indicates the binary mask resized to match the resolution of the input feature of $\text{EFA}_{C}^{a}$. We denote $E(\text{softmax}([\frac{QK^{\top}}{\sqrt{d}},\text{AP}_{a}(\mathbf{z}_{t}^{\text{cf}})]),V_{a})$ as $A_{a}^{\text{cf}}$ in Fig. 2 for simplicity.
@@
$\mathcal{L}_{\text{tot}}=\mathcal{L}_{\text{recon}}+\lambda_{1}\mathcal{L}_{\text{reg}}^{\text{cf}}+\lambda_{2}\mathcal{L}_{\text{reg}},$
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters to control relative importance between the losses. During training, only the parameters of $\text{AP}_{a}$ in $\text{EFA}_{C}^{a}$ are updated. Note that EFA is applied to the up-block layers with an input resolution of 16×16, as lower-resolution features capture higher-level semantic information. The choice of applied layers is discussed in Section 5.2. The method of extending EFA to handle multiple biases is provided in the Supplementary.
@@ 3.4 Inference
To mitigate the bias $C$ and ensure fair generation, we randomly sample an attribute $a_{i}$ with an equal probability from $\mathcal{A}_{C}$ and apply the corresponding $\text{EFA}_{C}^{a_{i}}$ to the original model. This approach enables users to easily control the attribute distribution by adjusting the sampling frequency, without additional training.
@@ 4.1 Experimental settings
Datasets. Following the previous studies [9, 10, 14], we mainly conducted experiments with gender and racial bias. Experiments with age bias are provided in Supplementary. The target attributes set of gender and racial bias is set as $\mathcal{A}_{\text{gender}}=\{\text{female},\text{male}\}$ and $\mathcal{A}_{\text{race}}=\{\text{White},\text{Black},\text{Asian},\text{Indian}\}$, respectively. For multiple biases, the set of target attributes is defined as $\mathcal{A}_{\text{gender}\times\text{race}}=\mathcal{A}_{\text{gender}}\times\mathcal{A}_{\text{race}}$. Note that gender is a complex concept that extends beyond a binary perspective. However, due to the lack of datasets and evaluation metrics that adequately represent diverse groups of gender, we adopt a binary perspective in this study. Further details of the dataset construction are included in the Supplementary.
@@
For non-target attribute preservation scores, we generate image pairs using the original Stable Diffusion (SD) [1] and our approach with the same random seed to evaluate how well non-target attributes are preserved while modifying target attributes. First, we segment the non-human regions in the original image and measure how well they are retained in the image generated by ours using Peak Signal-to-Noise Ratio (PSNR) and Learned Perceptual Image Patch Similarity (LPIPS) [28]. To assess overall image preservation, including contextual elements within human regions (e.g., a doctor’s stethoscope or white coat), we compute the DINO score, which measures the cosine similarity between the DINO [5] embeddings of images generated by SD and those generated by our method.
@@
Figure 5: Attention map visualization when mitigating gender bias using EFA. (a) and (b) show the results of applying $\text{EFA}_{\text{gender}}^{\text{female}}$ and $\text{EFA}_{\text{gender}}^{\text{male}}$, respectively. $<$sos$>$ and $<$eos$>$ indicate the start-of-sequence token and end-of-sequence token, respectively.
@@ 4.2 Experimental results
Bias mitigation and non-target attribute preservation. Table 1 presents the comparative experimental results of our method and baseline models using the WinoBias dataset. Overall, models struggle to maintain non-target attributes more in $\mathcal{T}_{\text{complex}}$ than $\mathcal{T}_{\text{basic}}$. Concept Algebra exhibits low performance in bias mitigation. Both UCE [10] and Finetuning Diffusion [25] modify model parameters directly, making it challenging to preserve the non-target attributes consistent with the original SD. Interpret Diffusion [14] demonstrates better bias mitigation performance compared to other baselines. However, it achieves low non-target attribute preservation scores due to its reliance on intermediate feature vectors to induce target attributes, applying uniform modifications without accounting for spatial context. In contrast, our method leverages spatial information to enhance target attributes at appropriate locations, thereby effectively preserving regions of non-target attributes.

### v2

@@ Abstract
Recent advancements in diffusion-based text-to-image (T2I) models have enabled the generation of high-quality and photorealistic images from text. However, they often exhibit societal biases related to gender, race, and socioeconomic status, thereby potentially reinforcing harmful stereotypes and shaping public perception in unintended ways. While existing bias mitigation methods demonstrate effectiveness, they often encounter attribute entanglement, where adjustments to attributes relevant to the bias (i.e., target attributes) unintentionally alter attributes unassociated with the bias (i.e., non-target attributes), causing undesirable distribution shifts. To address this challenge, we introduce Entanglement-Free Attention (EFA), a method that accurately incorporates target attributes (e.g., White, Black, and Asian) while preserving non-target attributes (e.g., background) during bias mitigation. At inference time, EFA randomly samples a target attribute with equal probability and adjusts the cross-attention in selected layers to incorporate the sampled attribute, achieving a fair distribution of target attributes. Extensive experiments demonstrate that EFA outperforms existing methods in mitigating bias while preserving non-target attributes, thereby maintaining the original model’s output distribution and generative capacity.
Figure 1: Generated images from the pretrained model (e.g., Stable Diffusion [1]), by previous debiasing methods (e.g., UCE [11] and Interpret Diffusion [15]), and ours. Previous debiasing methods often alter the non-target attributes (e.g., house) while modifying target bias attributes (e.g., racial attributes). In contrast, our approach preserves non-target attributes during bias mitigation (i.e., changing the racial attributes without modifying the house in the background), thereby preventing the introduction of unintended biases. The corresponding aligned images were generated using the same random seed.
@@ 1 Introduction
To address this issue, we propose Entanglement-Free Attention (EFA), a method that accurately incorporates target attributes while preserving non-target attributes. In this paper, we focus on human-centric bias, which plays a crucial role in shaping societal perceptions. Our EFA operates by leveraging predefined target attributes (e.g., White, Black, Asian, and Indian) that contribute to bias (e.g., race). To guide the model to focus on semantically relevant regions (e.g., face and body), EFA is trained to enhance target attributes in regions that exhibit high semantic relevance to the counterfactual prompts (i.e., prompts with alternative attributes other than the target), while human segmentation masks further constrain its effect to human regions. By employing both target and counterfactual prompts in training, EFA learns to dynamically adjust the enhancement strength based on the presence of the target attribute, applying stronger modifications when it is underrepresented. During inference, to ensure a fair distribution of target attributes in the generated output, one attribute is randomly sampled with equal probability. EFA then enhances the corresponding features in bias-relevant regions without relying on human-provided masks, effectively disentangling target and non-target attributes while preserving the original model’s generative capacity and contextual consistency.
@@ 2 Related works
Figure 2: Overview of our method. (a): EFA is applied to selected layers and trained to reconstruct the image $\mathbf{x}_{0}$ (a-1) when given $p_{a}$, the prompt containing the target attribute $a$, or (a-2) its counterfactual prompt, $p_{a}^{\text{cf}}$. We leverage a human segmentation mask $M$ to minimize the effect of EFA on the attention map of non-target attribute regions. (b): During the inference, our method generates outputs with a fair distribution of target attributes using the EFA of the randomly sampled target attribute.
@@
Existing studies on mitigating bias in T2I generative models have mainly followed two approaches. The first approach involves directly modifying the parameters of diffusion-based T2I models. For instance, Orgad et al. [19] and Gandikota et al. [11] proposed a method to adjust the weights of the cross-attention module to alleviate the bias. Similarly, Shen et al. [26] finetune the text encoder to alleviate face-related biases (e.g., gender, race). However, since finetuning is conducted on a significantly smaller dataset compared to the pretraining dataset, such approach may potentially degrade original output quality or diversity. Furthermore, the approach of Shen et al. [26] requires substantial computational costs for training the backbone model.
@@
Another line of research has explored inference-time steering approaches, where the pretrained model remains unchanged while adjustments are applied during inference. Friedrich et al. [10] proposed a bias mitigation framework that employs semantic guidance [4] to manipulate the target attributes by referring to registered information of a concept (e.g., nurse) and its bias (e.g., gender). However, this cannot mitigate bias for unregistered concepts, limiting its practicality. Chuang et al. [8] have developed a method to remove bias from text embeddings, and Kim et al. [13] trains the soft prompts for mitigating bias. Beyond employing the text embeddings, Shrestha et al. [27] introduced a technique that projects a reference image into text-compatible tokens, allowing the model to leverage the features of external images, such as gender or race for image generation. Additionally, Parihar et al. [20] and Li et al. [15] proposed methods to manipulate the h-space [14] of the model during inference. While these methods aim to mitigate bias without directly altering the backbone parameters, they often struggle to preserve the original context of the generated images.
@@
Recently, Zhou et al. [31] addressed association-engendered stereotypes arising from co-occurrence of multiple objects in an image. They rely on a CLIP model trained to predict biases from prompts, requiring prior knowledge of the biases to be mitigated. In contrast, our approach aims to address subtle biases induced by contextual variations beyond the predefined association bias definitions. For instance, if a particular social group is predominantly depicted in low-resource work environments, while another is shown in corporate office settings, such context variations can reinforce cognitive biases, subtly implying disparities in professional status. Furthermore, it is impractical to anticipate all possible forms of such bias in advance.
@@ 3.3 Training
Dataset construction. To train $\text{EFA}_{C}^{a_{i}}$ to represent a target attribute $a_{i}$ (e.g., female) while preserving non-target attributes, we create a dataset $\mathcal{D}_{a_{i}}$ that captures $a_{i}$ and provides pseudo supervision on the human regions that contain $a_{i}$. Specifically, we generate images using the original model with prompts explicitly describing $a_{i}$ (e.g., “a female person”). For each image, we extract a human segmentation mask $M$ using Grounded SAM2 [22, 17], assigning 1 to human regions and 0 elsewhere. Since this study focuses on human-centric bias, we utilize a human segmentation mask to serve as the basis for regulating the influence of $a_{i}$ during training. The resulting $\mathcal{D}_{a_{i}}$ consists of pairs of images and their segmentation masks.
Training EFA. At inference, the extent of the features that need to be adjusted by EFA to faithfully reflect $a_{i}$ may vary across inputs. For instance, with the prompt “nurse”, intermediate features may exhibit a high level of femininity, requiring minimal adjustment by $\text{EFA}_{\text{gender}}^{\text{female}}$. In contrast, intermediate features of the prompt “doctor” may have stronger masculinity, requiring enhanced representation of female by $\text{EFA}_{\text{gender}}^{\text{female}}$ to incorporate ‘female’ into the output. To handle this variability, $\text{EFA}_{C}^{a_{i}}$ is trained in two different scenarios : (1) when intermediate features contain the attribute $a_{i}$ (e.g., female), and (2) when intermediate features contain its counterfactual attribute $a_{i}^{\text{cf}}\in\mathcal{A}_{C}\setminus\{a_{i}\}$ (e.g., male). We denote $a_{i}$ as $a$ throughout this section for simplicity.
@@
$\mathcal{L}_{\text{reg}}(\mathbf{z}_{t},W)={\lVert W\odot A_{a}(\mathbf{z}_{t})\rVert_{1}},$
where $\mathbf{z}_{t}=\phi(\textbf{x}_{t},p_{a})$, $W$ is a spatial mask used to constrain the effect of the loss, and $\odot$ represents the element-wise multiplication. Also, $A_{a}(\mathbf{z}_{t})$ denotes the attention values (after softmax) associated with the value vectors $V_{a}$ given $\mathbf{z}_{t}$. More precisely, when $A(\mathbf{z}_{t})=\text{softmax}([\frac{QK^{\top}}{\sqrt{d}},\text{AP}_{a}(\mathbf{z}_{t})])$, we define $A_{a}(\mathbf{z}_{t})$ as the submatrix of $A(\mathbf{z}_{t})$ obtained by selecting values indexed by $I_{V_{a}}$, where $I_{V_{a}}$ denotes the set of indices corresponding to $V_{a}$.
In this scenario, we set $W=\mathbf{1}$, a mask of ones, to apply the loss over all spatial positions. This regularization loss minimizes unnecessary modifications to the original cross-attention output introduced by $\text{EFA}_{C}^{a}$. Hereafter, we denote $\mathcal{L}_{\text{reg}}(\mathbf{z}_{t},\mathbf{1})$ as $\mathcal{L}_{\text{reg}}^{\text{trg}}$ for brevity.
@@
where $\epsilon_{\theta}$ denotes the model with $\text{EFA}_{C}^{a}$. Through the reconstruction process, $\text{EFA}_{C}^{a}$ is encouraged to enhance $a$ in regions where cross-attention assigns high similarity to $a^{\text{cf}}$, reflecting their semantic relevance. Using human masks ensures that $a$ is accurately reconstructed in human regions while mitigating the influence of $a^{\text{cf}}$.
Although $\mathcal{L}_{\text{reg}}^{\text{trg}}$ reduces unintended modifications, it does not explicitly confine its impact on human regions, potentially affecting non-target attributes. To address this, we impose an additional L1 regularization on the attention values corresponding to $V_{a}$, applied outside the human region, as follows:
$\mathcal{L}_{\text{reg}}^{\text{cf}}=\mathcal{L}_{\text{reg}}(\mathbf{z}_{t}^{\text{cf}},\mathbf{1}-\bar{M}),$
where $\mathbf{z}_{t}^{\text{cf}}=\phi(\textbf{x}_{t},p_{a}^{\text{cf}})$, and $\bar{M}$ indicates the binary mask resized to match the resolution of the input feature of $\text{EFA}_{C}^{a}$.
@@
$\mathcal{L}_{\text{tot}}=\mathcal{L}_{\text{recon}}+\mathcal{L}_{\text{reg}}^{*},$
where $\mathcal{L}_{\text{reg}}^{*}=\lambda_{1}\mathcal{L}_{\text{reg}}^{\text{trg}}+\lambda_{2}\mathcal{L}_{\text{reg}}^{\text{cf}}$ and $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters to control the relative importance between the losses. During training, only the parameters of $\text{AP}_{a}$ in $\text{EFA}_{C}^{a}$ are updated. Note that EFA is applied to the up-block layers with an input resolution of 16×16, as lower-resolution features capture higher-level semantic information. The choice of applied layers is discussed in Section 5.2. The method of extending EFA to handle multiple biases is provided in the Supplementary.
@@ 3.4 Inference
To mitigate the bias $C$ and ensure fair generation, we randomly sample an attribute $a_{i}$ with an equal probability from $\mathcal{A}_{C}$ and apply the corresponding $\text{EFA}_{C}^{a_{i}}$ to the original model. This approach enables users to easily control the attribute distribution by adjusting the sampling frequency, without additional training. Note that our method requires no masks at inference and generates debiased images directly from noise, offering a practical and efficient solution.
@@ 4.1 Experimental settings
Datasets. Following the previous studies [10, 11, 15], we mainly conducted experiments with gender and racial bias. Experiments with age bias are provided in Supplementary. The target attributes set of gender and racial bias is set as $\mathcal{A}_{\text{gender}}=\{\text{female},\text{male}\}$ and $\mathcal{A}_{\text{race}}=\{\text{White},\text{Black},\text{Asian},\text{Indian}\}$, respectively. For multiple biases, the set of target attributes is defined as $\mathcal{A}_{\text{gender}\times\text{race}}=\mathcal{A}_{\text{gender}}\times\mathcal{A}_{\text{race}}$. We adopt a binary gender perspective to evaluate our method, following prior works [11, 15], due to the limitations of existing datasets and evaluation protocols. Further details of the dataset construction are included in the Supplementary.
@@
For non-target attribute preservation scores, we generate image pairs using the original Stable Diffusion (SD) [1] and our approach with the same random seed to evaluate how well non-target attributes are preserved while modifying target attributes. First, we segment the non-human regions in the original image and measure how well they are retained in the image generated by ours using Peak Signal-to-Noise Ratio (PSNR) and Learned Perceptual Image Patch Similarity (LPIPS) [29]. To assess overall image preservation, including contextual elements within human regions (e.g., a doctor’s stethoscope or white coat), we compute the DINO score, employing DINO’s ability to capture low-level and semantically meaningful visual details. The score measures the cosine similarity between the DINO [5] embeddings of images generated by SD and our method.
@@
Figure 5: Attention map visualization when mitigating gender bias using EFA. (a) and (b) show the results of applying $\text{EFA}_{\text{gender}}^{\text{female}}$ and $\text{EFA}_{\text{gender}}^{\text{male}}$, respectively. $<$sos$>$ and $<$eos$>$ denote special tokens that mark the beginning and end of a text sequence, respectively.
@@ 4.2 Experimental results
Bias mitigation and non-target attribute preservation. Table 1 presents the comparative experimental results of our method and baseline models using the WinoBias dataset. Overall, models struggle to maintain non-target attributes more in $\mathcal{T}_{\text{complex}}$ than $\mathcal{T}_{\text{basic}}$. Concept Algebra exhibits low performance in bias mitigation. Both UCE [11] and Finetuning Diffusion [26] modify model parameters directly, making it challenging to preserve the non-target attributes consistent with the original SD. Interpret Diffusion [15] demonstrates better bias mitigation performance compared to other baselines. However, it exhibits low non-target attribute preservation scores due to its uniform application of fixed semantic vectors without spatial awareness. In contrast, our input-aware architecture adaptively enhances target attributes based on spatial context, enabling effective preservation of non-target attributes.